# The metaplectic semigroup and its applications to time-frequency analysis and evolution operators

### v1

@@ 1.4.3. Schrödinger representation
(25) $f\in\mathrm{span}\big\{\text{$\rho(z;\tau)e^{-\pi Mx\cdot x}$ : $z\in{\mathbb{R}^{2d}}$, $\tau\in\mathbb{R}$, and $M=M^{\top}$ is positive-definite}\big\}.$
We address Section 4.3 for the details. A similar result appears in [58, Section 5], although the link between the metaplectic operator and its symplectic projection is not discussed there. Here, we present an explicit version of this relation. Theorem 1.5 is used in Section 6 to characterize covariant metaplectic Wigner distributions.
@@ 1.4.4. Intertwining with complex conjugation and its applications
For operators in $\operatorname*{Mp}(d,\mathbb{R})$, the following result was proven in [13, Appendix A].
@@
From the rest of this work, the operator $T$ of Theorem 1.7 will be denoted by $\widehat{\widetilde{S}}$, consistently with Proposition 1.6. The reader may observe that in the complex case it is not true anymore that $\widetilde{S}=RSR$, so complex conjugation alone cannot be included consistently among metaplectic operators in the complex framework. The subsemigroup of metaplectic operators so that $\widehat{\widetilde{S}}=\widehat{S}$, that are precisely those commuting with complex conjugation, are then characterized in Theorem 5.9. Remarkably, its proof requires a preliminary result on the structure of $\operatorname*{Sp}(d,\mathbb{R})$, see Lemma 5.8.
@@ 1.4.5. A new metaplectic setting for time-frequency representations
By allowing $\widehat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{C})$ in Definition 1.1, we obtain a generalization of metaplectic Wigner distributions of [18], that we present in Section 6. We thereby transport the techniques resulted advantageous in the framework of the metaplectic group $\operatorname*{Mp}(2d,\mathbb{R})$ to the metaplectic semigroup $\operatorname*{Mp}_{+}(2d,\mathbb{C})$. Importantly, we characterize the following classes of metaplectic Wigner distributions in terms of both the symplectic projection $\mathcal{A}$ of the metaplectic operator $\widehat{\mathcal{A}}$ defining $W_{\mathcal{A}}$, and provide the explicit expressions of the time-frequency representations satisfying these properties.
@@
Finally, the (complex) metaplectic theory that we study applies also to evolution equations with quadratic non-selfadjoint Hamiltonian. In this direction, we consider Cauchy problems of the form
@@
complex quadratic form associated with a matrix $\mathcal{Q}\in\mathbb{C}^{2d\times 2d}$ satisfying $\Re(\mathcal{Q})\geq 0$. For problems of this form it is well known that there exists a solution operator, that we usually call propagator,
@@
and moreover it is a one-parameter semigroup of metaplectic operators $e^{-t\mathrm{Op}^{\mathrm{w}}(a)}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ (cf. [41]). Hence, in Section 7 we study the evolution of the microlocal concentration of $\mathcal{A}$-Wigner distributions, with $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})$ covariant, induced by (35). In such a section, relying on the evolution of a mixed heat–Schrödinger type induced by harmonic oscillators, we prove boundedness results on modulation spaces and propagation of micro-singularities, we address to Theorem 7.10 for a precise statement.
@@ 2. Preliminaries and notation
We denote by $xy=x\cdot y$ the standard inner product of $x,y\in\mathbb{R}^{d}$. For a $d\times d$ complex matrix $S\in\mathbb{C}^{d\times d}$, we write $\Re(S)$ and $\Im(S)$ to denote its real and imaginary parts, respectively, while $\overline{S}$ denotes its complex conjugate. We denote by $\operatorname*{GL}(d,\mathbb{R})$, respectively $\operatorname*{GL}(d,\mathbb{C})$, the group of $d\times d$ real, respectively complex, invertible matrices. Analogously, $\operatorname*{Sym}(d,\mathbb{R})$ and $\operatorname*{Sym}(d,\mathbb{C})$, denote the class of symmetric $d\times d$ real and complex matrices, i.e., satisfying $S^{\top}=S$. If $S\in\operatorname*{Sym}(d,\mathbb{R})$, we write $S\geq 0$, respectively $S>0$, if $S$ is positive semi-definite, respectively positive-definite. We also write $S\leq 0$ and $S<0$ if $S\geq 0$ and S>0-S>0, respectively. We denote by $\Sigma_{\geq 0}(d)$ the (closure of the) Siegel upper-half space, i.e., the set of complex symmetric matrices $S$ with $\Im(S)\geq 0$. We denote by $I_{d}$ and $O_{d}$ the $d\times d$ identity and null matrices, respectively. We also write $O_{m\times n}$ for the $m\times n$ null matrix. If $d$ is clear from the context, it will be omitted. If $S_{1},S_{2}$ are $2d\times 2d$ matrices, with
@@
The translation and the modulation of a function $f\in L^{2}(\mathbb{R}^{d})$ are $T_{x}f(t)=f(t-x)$ and $M_{\xi}f(t)=e^{2\pi i\xi\cdot t}f(t)$, respectively. For $M>0$, we write $\varphi_{M}(t)=e^{-\pi Mt\cdot t}$ for the Gaussian with covariance matrix $M$.
In what follows, we present briefly the background needed for our discussion. For a more exhaustive treatment of these topics, the interested reader may consult [17, 27, 36] and the reference therein.
@@ 2.1. Pseudo-inverses and Schur complements
Schur complements are used in this work to study semi-definiteness of symmetric matrices. Let us recall the definition of Moore-Penrose inverse of a real-valued matrix $M$. There exist $U,V$ orthogonal and $\Sigma$ diagonal such that $M=U\Sigma V^{\top}$ and the diagonal elements of $\Sigma$ are the singular values of $M$, that are the eigenvalues of $(M^{\top}M)^{1/2}$. Without loss of generality, we may assume that $\Sigma=\operatorname*{diag}(\sigma_{1},\ldots,\sigma_{d})$, with $\sigma_{1}\geq\ldots\geq\sigma_{r}\geq 0$, $r=\mathrm{rank}(M)$. The Moore-Penrose inverse of $M$ is $M^{+}=V\Sigma^{+}U^{\top}$, where $\Sigma^{+}=\operatorname*{diag}(\lambda_{1},\ldots,\lambda_{d})$, $\lambda_{j}=\sigma_{j}^{-1}$ for $1\leq j\leq r$ and $\sigma_{j}=0$ if $r<j\leq d$. We recall that $M^{+}M$ is the orthogonal projection onto $\ker(M)^{\perp}$ and $MM^{+}$ is the orthogonal projection onto $R(M)$. Both $M^{+}M$ and $MM^{+}$ are symmetric and positive semi-definite. Finally, we recall that $(M^{+})^{\top}=(M^{\top})^{+}$.
@@
Since $AA^{+}$ is the orthogonal projection onto $R(A)$, the reader may observe that $(I-AA^{+})B=O_{d}$ is then equivalent to $R(B)\subseteq R(A)$. Similarly, $(I-CC^{+})B^{\top}=O_{d}$ is equivalent to $R(B^{\top})\subseteq R(C)$.
@@ 2.2. Gaussian integrals
Let $M\in\operatorname*{Sym}(d,\mathbb{C})\cap\operatorname*{GL}(d,\mathbb{C})$ be such that $\Re(M)\geq 0$. Then, for every $z\in\mathbb{C}^{d}$,
@@
where $\det(M)^{-1/2}$ is chosen so that $\det(M)^{-1/2}>0$ if $M$ is real.
The integral in (40) must be interpreted as the distributional sense if $\Re(M)$ is not invertible.
@@ 2.3. The real and complex symplectic groups
We refer the reader to [17, 27, 49] as exhaustive sources. We denote by
@@
(42) $\sigma(x,\xi)=Jz\cdot w,\qquad z,w\in{\mathbb{R}^{2d}}.$
@@
We recall that every symplectic matrix $U\in\operatorname*{Sp}(d,\mathbb{R})$ admits a singular values decomposition of symplectic matrices. Specifically, there exist $V,W\in\operatorname*{Sp}(d,\mathbb{R})$ orthogonal and $\delta\in\operatorname*{Sp}(d,\mathbb{R})$ diagonal such that $U=W\delta V^{\top}$. The diagonal elements of $\delta$ are the singular values of $U$. Recall that if $\lambda$ is a singular value of $U$, then also $1/\lambda$ is a singular value of $U$. Observe that this also implies that $U$ and $U^{-1}$ have the same singular values, and that the largest singular value of $U$ is $\geq 1$. We refer to [55] for the details. Moreover, by imposing $V^{-1}=V^{\top}$, it is easy to see that $V\in\operatorname*{Sp}(d,\mathbb{R})$ is orthogonal if and only if
@@
Now, we discuss the generators of the symplectic group. For $E\in\operatorname*{GL}(d,\mathbb{C})$ and $Q\in\operatorname*{Sym}(d,\mathbb{C})$, we set
@@
The symplectic group $\operatorname*{Sp}(d,\mathbb{C})$ is path-connected [49, Theorem 4.7]. Hence, the very same argument proving [27, Proposition 4.9] shows that $\operatorname*{Sp}(d,\mathbb{C})$ is generated by the complex analogues of the real generators. Precisely,
@@ 2.4. The metaplectic group
[17, 27] For $x,\xi\in\mathbb{R}^{d}$ and $\tau\in\mathbb{R}$, we consider the operator
@@
Then, $\rho$ is a unitary irreducible representation of the Heisenberg group on $L^{2}(\mathbb{R}^{d})$, and it is called the Schrödinger representation. For every $U\in\operatorname*{Sp}(d,\mathbb{R})$, the representation $\rho_{U}(x,\xi;\tau)=\rho(U(x,\xi);\tau)$ is another irreducible representation of the Heisenberg group, with $\rho_{U}(0;\tau)=e^{2\pi i\tau}\operatorname*{id}_{L^{2}}$. By the Stone-von Neumann theorem, $\rho$ and $\rho_{U}$ are unitarily equivalent, i.e., there exists $\widehat{U}:L^{2}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$ unitary so that
@@
The metaplectic group is generated by
@@
and $m\in\mathbb{Z}$ is the argument of $\det(E)$ (Maslov index). Moreover, $\pi^{Mp}(\mathcal{F})=J$, $\pi^{Mp}(\mathfrak{p}_{Q})=V_{Q}$ and $\pi^{Mp}(\mathfrak{T}_{E})=\mathcal{D}_{E}$, these matrices being defined as in (41), (51) and (52), respectively.
When it does not cause confusion, we shall omit phase factors in (55) and (57).
@@
We will use the following result, see [20, Proposition 105] with $\hbar=1/2\pi$.
@@
Let $\widehat{U}\in\operatorname*{Mp}(d,\mathbb{R})$ be orthogonal with projection as in (47). Let $\varphi(x)=e^{-\pi|x|^{2}}$. Then,
@@ 2.6. Pseudodifferential operators
[42, 47] If $f,g\in\mathcal{S}(\mathbb{R}^{d})$, then $W(f,g)\in\mathcal{S}({\mathbb{R}^{2d}})$. If $a\in\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$, we may consider the Weyl-quantized pseudodifferential operator (or, Weyl operator) with symbol $a$, that is the linear operator continuous from $\mathcal{S}(\mathbb{R}^{d})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ defined by
@@
where the integral is interpreted in a distributional sense, see [17, Chapter 4] for the details.
We conclude this subsection by recalling the metaplectic invariance of Weyl Hörmander calculus, stated as follows (see [47, Theorem 2.1.1]).
@@ 2.7. Modulation spaces
$M^{2}(\mathbb{R}^{d})=L^{2}(\mathbb{R}^{d})$ with equivalence of norms. Moreover, for $s\in\mathbb{R}$, the Shubin-Sobolev space
(71) $Q_{s}(\mathbb{R}^{d})=\{f\in\mathcal{S}^{\prime}(\mathbb{R}^{d}):\lVert f\rVert_{Q_{s}}^{2}=\sum_{|\alpha|+|\beta|<\infty}\lVert x^{\alpha}D^{\beta}_{x}f\rVert_{2}<\infty\},\qquad s\in\mathbb{R},$
coincides with $M^{2}_{v_{s}}(\mathbb{R}^{d})$, with equivalence of norms, cf. [17, Section 4.4.3.1].
@@
(72) $S^{0}_{0,0}(\mathbb{R}^{d})=\{f\in\mathcal{C}^{\infty}(\mathbb{R}^{d}):D^{\alpha}f\in L^{\infty}(\mathbb{R}^{d}),\forall\alpha\in\mathbb{N}^{d}\}=\bigcap_{s\geq 0}M^{\infty,1}_{1\otimes v_{s}}(\mathbb{R}^{d}).$
@@
The boundedness of metaplectic operators in $\operatorname*{Mp}(d,\mathbb{R})$ on modulation spaces has been studied in [28]. A more quantitative approach, more similar to ours, was used in in [16, Corollary 3.5] for the unweighted $M^{p}$-spaces. Here, we obtain estimates for $\lVert\widehat{U}\rVert_{op}$ in terms of the singular values of the projection $U$, for $M^{p,q}_{v_{s}}$-spaces.
@@
where $c=c(g,s,d,p,q)>0$, being $g$ the window chosen to compute the $M^{p,q}_{v_{s}}$-norm, and we understand $1/\infty=1$ and $|\det(A)|^{1/p-1/q}=1$ if $p=q$. Moreover,
(75) $\lVert\widehat{U}f\rVert_{M^{p,q}_{v_{s}}}\geq\frac{1}{c}|\det(A)|^{1/p-1/q}{\sigma_{\mathrm{max}}(U)^{-2s}}\prod_{j=1}^{d}\Big(\frac{\sigma_{j}}{1+\sigma_{j}^{2}}\Big)^{1/2}\lVert f\rVert_{M^{p,q}_{v_{s}}}.$
@@
The statements concerning the isomorphisms were proven in [28]. We just need to check the upper bounds (74) and (75). Let $f\in M^{p,q}_{v_{s}}(\mathbb{R}^{d})$ and $g\in\mathcal{S}(\mathbb{R}^{d})\setminus\{0\}$ be a window function. Similarly to (5),
@@
where we interpret $|\det(A)|^{1/p-1/q}=1$ if $p=q$. Since $v_{s}(Uz)\leq\sigma_{\mathrm{max}}(U)^{s}v_{s}(z)$, where $\sigma_{\mathrm{max}}(U)$ is the largest singular value of $U$, we obtain
@@
Next, we replace the window $\widehat{U}^{-1}g$ with $g$ by resorting to the standard convolution inequality in [17, Lemma 1.2.29] to obtain
@@
By Young’s inequality for weighted mixed norm spaces, see e.g. [17],
(88) $\lVert|V_{g}f|\ast|V_{\widehat{U}^{-1}g}\widehat{U}^{-1}g|\rVert_{L^{p,q}_{v_{s}}}$ $\leq c^{\prime}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}}\lVert V_{\widehat{U}^{-1}g}\widehat{U}^{-1}g\rVert_{L^{1}_{v_{s}}}=c^{\prime}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}}\lVert V_{g}g(U\cdot)\rVert_{L^{1}_{v_{s}}}$
(89) $\leq c^{\prime}\sigma_{\mathrm{max}}(U^{-1})^{s}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}}\lVert V_{g}g\rVert_{L^{1}_{v_{s}}}=c^{\prime}\sigma_{\mathrm{max}}(U)^{s}\lVert V_{g}g\rVert_{L^{1}_{v_{s}}}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}},$
@@
On the other hand, the computation of $V_{g}g$ is standard, see [17, Lemma 1.3.36],
(95) $|V_{g}g(x,\xi)|=2^{-d/2}e^{-2\pi|x|^{2}}e^{-\pi|\xi|^{2}/2},\qquad x,\xi\in\mathbb{R}^{d},$
so, by using that $v_{s}(x/\sqrt{2},\sqrt{2}\xi)\leq 2^{s/2}v_{s}(x,\xi)$,
(96) $\lVert V_{g}g\rVert_{L^{1}_{v_{s}}}=2^{(s-d)/2}\int_{{\mathbb{R}^{2d}}}e^{-\pi|z|^{2}}v_{s}(z)dz.$

### v2

@@ 1.4.3. Schrödinger representation
(25) $f\in\mathrm{span}\big\{\text{$\rho(z;\tau)e^{-\pi Mx\cdot x}$ : $z\in{\mathbb{R}^{2d}}$, $\tau\in\mathbb{R}$, and $M=M^{\top}$ real and positive-definite}\big\}.$
We address Section 4.3 for the details. A similar result appears in [56, Section 5], although the link between the metaplectic operator and its symplectic projection is not discussed there. Here, we present an explicit version of this relation. Theorem 1.5 is used in Section 6 to characterize covariant metaplectic Wigner distributions.
@@ 1.4.4. Intertwining with complex conjugation and its applications
For operators in $\operatorname*{Mp}(d,\mathbb{R})$, the following result was proven in [12, Appendix A].
@@
For the rest of this work, the operator $T$ of Theorem 1.7 will be denoted by $\widehat{\widetilde{S}}$, consistently with Proposition 1.6. The reader may observe that in the complex case it is not true anymore that $\widetilde{S}=RSR$, so complex conjugation alone cannot be included consistently among metaplectic operators in the complex framework. The subsemigroup of metaplectic operators so that $\widehat{\widetilde{S}}=\widehat{S}$, that are precisely those commuting with complex conjugation, are then characterized in Theorem 5.10. Remarkably, its proof requires a preliminary result on the structure of $\operatorname*{Sp}(d,\mathbb{R})$, see Lemma 5.9.
@@ 1.4.5. A new metaplectic setting for time-frequency representations
By allowing $\widehat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{C})$ in Definition 1.1, we obtain a generalization of metaplectic Wigner distributions of [17], that we present in Section 6. We thereby transport the techniques resulted advantageous in the framework of the metaplectic group $\operatorname*{Mp}(2d,\mathbb{R})$ to the metaplectic semigroup $\operatorname*{Mp}_{+}(2d,\mathbb{C})$. Importantly, we characterize the following classes of metaplectic Wigner distributions in terms of the symplectic projection $\mathcal{A}$ of the metaplectic operator $\widehat{\mathcal{A}}$ defining $W_{\mathcal{A}}$, and provide the explicit expressions of the time-frequency representations satisfying these properties.
@@
Finally, the (complex) metaplectic theory that we study applies to evolution equations with quadratic non-selfadjoint Hamiltonian. In this direction, we consider Cauchy problems of the form
@@
complex quadratic form associated with a matrix $\mathcal{Q}\in\mathbb{C}^{2d\times 2d}$ satisfying $\Re(\mathcal{Q})\geq 0$. As aformentioned, the propagator
@@
is a one-parameter semigroup of metaplectic operators $e^{-t\mathrm{Op}^{\mathrm{w}}(a)}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ (cf. [39]). Hence, in Section 7 we study the evolution of the microlocal concentration of $\mathcal{A}$-Wigner distributions, with $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})$ covariant, induced by (35). In such a section, relying on the evolution of a mixed heat–Schrödinger type induced by harmonic oscillators, we prove boundedness results on modulation spaces and propagation of micro-singularities, we address Section 7.
@@ 2. Preliminaries and notation
We denote by $xy=x\cdot y$ the standard inner product of $x,y\in\mathbb{R}^{d}$. For a $d\times d$ complex matrix $S\in\mathbb{C}^{d\times d}$, we write $\Re(S)$ and $\Im(S)$ to denote its real and imaginary parts, respectively, while $\overline{S}$ denotes its complex conjugate. We denote by $\operatorname*{GL}(d,\mathbb{R})$, respectively $\operatorname*{GL}(d,\mathbb{C})$, the group of $d\times d$ real, respectively complex, invertible matrices. Analogously, $\operatorname*{Sym}(d,\mathbb{R})$ and $\operatorname*{Sym}(d,\mathbb{C})$, denote the class of symmetric $d\times d$ real and complex matrices, i.e., satisfying $S^{\top}=S$. If $S\in\operatorname*{Sym}(d,\mathbb{R})$, we write $S\geq 0$, respectively $S>0$, if $S$ is positive semi-definite, respectively positive-definite. We also write $S\leq 0$ and $S<0$ if $S\geq 0$ and S>0-S>0, respectively. We denote by $\Sigma_{\geq 0}(d)$ the (closure of the) Siegel upper-half space, i.e., the set of complex symmetric matrices $S$ with $\Im(S)\geq 0$. We denote by $I_{d}$ and $O_{d}$ the $d\times d$ identity and null matrices, respectively. We also write $O_{m\times n}$ for the $m\times n$ null matrix. If $S_{1},S_{2}$ are $2d\times 2d$ matrices, with
@@
The translation and the modulation of a function $f\in L^{2}(\mathbb{R}^{d})$ are $T_{x}f(y)=f(y-x)$ and $M_{\xi}f(t)=e^{2\pi i\xi\cdot y}f(y)$, respectively. We also write $\pi(x,\xi)f=M_{\xi}T_{x}$. For $M>0$, we write $\varphi_{M}(t)=e^{-\pi Mt\cdot t}$ for the Gaussian with covariance matrix $M$.
In what follows, we present briefly the background needed for our discussion. We will mainly follow the notation and the terminology of [16] for time-frequency analysis and of [40, 45] for microlocal analysis. Equally valid sources are [22, 19, 27, 33].
@@ 2.1. Pseudo-inverses and Schur complements
Schur complements are used in this work to study semi-definiteness of symmetric matrices. Let us recall the definition of Moore-Penrose inverse of a real-valued matrix $M$. There exist $U,V$ orthogonal and $\Sigma$ diagonal such that $M=U\Sigma V^{\top}$ (singular value decomposition) and the diagonal elements of $\Sigma$ are the singular values of $M$, that are the eigenvalues of $(M^{\top}M)^{1/2}$. Without loss of generality, we may assume that $\Sigma=\operatorname*{diag}(\sigma_{1},\ldots,\sigma_{d})$, with $\sigma_{1}\geq\ldots\geq\sigma_{r}\geq 0$, $r=\mathrm{rank}(M)$. The Moore-Penrose inverse of $M$ is $M^{+}=V\Sigma^{+}U^{\top}$, where $\Sigma^{+}=\operatorname*{diag}(\lambda_{1},\ldots,\lambda_{d})$, $\lambda_{j}=\sigma_{j}^{-1}$ for $1\leq j\leq r$ and $\sigma_{j}=0$ if $r<j\leq d$. We recall that $M^{+}M$ is the orthogonal projection onto $\ker(M)^{\perp}$ and $MM^{+}$ is the orthogonal projection onto $\mathrm{range}(M)$. Both $M^{+}M$ and $MM^{+}$ are symmetric and positive semi-definite. Finally, we recall that $(M^{+})^{\top}=(M^{\top})^{+}$.
@@
Since $AA^{+}$ is the orthogonal projection onto $\mathrm{range}(A)$, the reader may observe that $(I-AA^{+})B=O_{d}$ is then equivalent to $\mathrm{range}(B)\subseteq\mathrm{range}(A)$. Similarly, $(I-CC^{+})B^{\top}=O_{d}$ is equivalent to $\mathrm{range}(B^{\top})\subseteq\mathrm{range}(C)$.
@@ 2.2. Gaussian integrals
Let $M\in\operatorname*{Sym}(d,\mathbb{C})\cap\operatorname*{GL}(d,\mathbb{C})$ be such that $\Re(M)>0$. Then, for every $z\in\mathbb{C}^{d}$,
@@
where $\det(M)^{-1/2}$ is chosen so that $\det(M)^{-1/2}>0$ if $M$ is real. Moreover, if $\Re(M)\geq 0$, (40) still holds in the sense of distribution.
@@ 2.3. The real and complex symplectic groups
We denote by
@@
(42) $\sigma(z,w)=J_{d}z\cdot w,\qquad z,w\in{\mathbb{R}^{2d}}.$
@@
We recall that every symplectic matrix $U\in\operatorname*{Sp}(d,\mathbb{R})$ admits a singular values decomposition with symplectic factors. Specifically, there exist $V,W\in\operatorname*{Sp}(d,\mathbb{R})$ orthogonal and $\delta\in\operatorname*{Sp}(d,\mathbb{R})$ diagonal such that $U=W\delta V^{\top}$. The diagonal elements of $\delta$ are the singular values of $U$. Recall that if $\lambda$ is a singular value of $U$, then also $1/\lambda$ is a singular value of $U$. Observe that this also implies that $U$ and $U^{-1}$ have the same singular values, and that the largest singular value of $U$ is $\geq 1$. We refer to [53] for the details. Moreover, by imposing $V^{-1}=V^{\top}$, it is easy to see that $V\in\operatorname*{Sp}(d,\mathbb{R})$ is orthogonal if and only if
@@
Now, we discuss the generators of the real and the complex symplectic groups. For $E\in\operatorname*{GL}(d,\mathbb{C})$ and $Q\in\operatorname*{Sym}(d,\mathbb{C})$, we set
@@
The symplectic group $\operatorname*{Sp}(d,\mathbb{C})$ is path-connected [47, Theorem 4.7]. Hence, the very same argument proving [27, Proposition 4.9] shows that $\operatorname*{Sp}(d,\mathbb{C})$ is generated by the complex analogues of the real generators. Precisely,
@@ 2.4. The metaplectic group
For $x,\xi\in\mathbb{R}^{d}$ and $\tau\in\mathbb{R}$, we consider the operator
@@
Then, $\rho$ is a unitary irreducible representation of the Heisenberg group on $L^{2}(\mathbb{R}^{d})$, and it is called the Schrödinger representation. For every $U\in\operatorname*{Sp}(d,\mathbb{R})$, the representation $\rho_{U}(x,\xi;\tau)=\rho(U(x,\xi);\tau)$ is another irreducible representation of the Heisenberg group, with $\rho_{U}(0;\tau)=e^{2\pi i\tau}\operatorname*{id}_{L^{2}}$. By Stone-von Neumann theorem, $\rho$ and $\rho_{U}$ are unitarily equivalent, i.e., there exists $\widehat{U}:L^{2}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$ unitary so that
@@
The metaplectic group $\operatorname*{Mp}(d,\mathbb{R})$ is generated by
@@
and $m\in\mathbb{Z}$ is the argument of $\det(E)$ (Maslov index). Moreover, $\pi^{Mp}(\mathcal{F})=J$, $\pi^{Mp}(\mathfrak{p}_{Q})=V_{Q}$ and $\pi^{Mp}(\mathfrak{T}_{E})=\mathcal{D}_{E}$, being these matrices defined as in (41), (51) and (52), respectively.
When it does not cause confusion, we shall omit the phase factors in (55) and (57).
@@
We will use the following result, see [19, Proposition 105] with $\hbar=1/2\pi$.
@@
Let $\widehat{U}\in\operatorname*{Mp}(d,\mathbb{R})$ with orthogonal projection as in (47). Let $\varphi(x)=e^{-\pi|x|^{2}}$. Then,
@@ 2.6. Pseudodifferential operators
If $f,g\in\mathcal{S}(\mathbb{R}^{d})$, then $W(f,g)\in\mathcal{S}({\mathbb{R}^{2d}})$. If $a\in\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$, we may consider the Weyl-quantized pseudodifferential operator (or, Weyl operator) with symbol $a$, that is the linear operator continuous from $\mathcal{S}(\mathbb{R}^{d})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ defined by
@@
where the integral is interpreted in a distributional sense, see [16, Chapter 4] for the details.
We conclude this subsection by recalling the metaplectic invariance of Weyl Hörmander calculus, stated as follows (see [45, Theorem 2.1.1]).
@@ 2.7. Modulation spaces
$M^{2}(\mathbb{R}^{d})=L^{2}(\mathbb{R}^{d})$ with equivalence of norms. Moreover, for $s\in\mathbb{R}$, the modulation space $M^{2}_{v_{s}}(\mathbb{R}^{d})$ coincides with the Shubin-Sobolev space which is defined as
(71) $Q_{s}(\mathbb{R}^{d})=\Big\{f\in\mathcal{S}^{\prime}(\mathbb{R}^{d});\ \Lambda^{s}f\in L^{2}(\mathbb{R}^{d})\Big\},\qquad s\in\mathbb{R},$
with norm,
$\|u\|_{Q_{s}}:=\|\Lambda^{s}u\|_{2},\quad u\in Q_{s}(\mathbb{R}^{d}),$
for $\Lambda^{s}=(1+|x|^{2}-\Delta)^{s/2}$ defined through the functional calculus, see [16, Section 4.4.3.1] and [34]. More simplicity, for $k\in\mathbb{N}$, the Shubin-Sobolev space of order $k$ can be equivalently defined as
$Q_{k}(\mathbb{R}^{d})=\{u\in L^{2}(\mathbb{R}^{d});\quad x^{\alpha}\partial_{x}^{\beta}u\in L^{2}(\mathbb{R}^{d}),\ \forall\lvert\alpha\rvert+\lvert\beta\rvert\leq k\}$
and
$\|u\|_{Q_{k}}=\Bigl(\sum_{\lvert\alpha\rvert+\lvert\beta\rvert\leq k}\|x^{\alpha}\partial_{x}^{\beta}u\|^{2}_{2}\Bigr)^{2}.$
@@
(72) $S^{0}_{0,0}(\mathbb{R}^{d})=\bigcap_{s\geq 0}M^{\infty,1}_{1\otimes v_{s}}(\mathbb{R}^{d}).$
@@
The boundedness of metaplectic operators in $\operatorname*{Mp}(d,\mathbb{R})$ on modulation spaces has been studied in [28]. A more quantitative approach, more similar to ours, was used in in [15, Corollary 3.5] for the unweighted $M^{p}$-spaces. Here, we obtain estimates for $\lVert\widehat{U}\rVert_{op}$ in terms of the singular values of the projection $U$, for $M^{p,q}_{v_{s}}$-spaces.
@@
where $c=c(g,s,d,p,q)>0$, being $g$ the window chosen to compute the $M^{p,q}_{v_{s}}$-norm, and we understand $1/\infty=1$ and $|\det(A)|^{1/p-1/q}=1$ if $p=q$ (also when $A=O_{d}$).
@@
The statements concerning the isomorphisms were proven in [28]. We need to check the upper bound (74). Let $f\in M^{p,q}_{v_{s}}(\mathbb{R}^{d})$ and $g\in\mathcal{S}(\mathbb{R}^{d})\setminus\{0\}$ be a window function. Similarly to (5),
@@
where we interpret $|\det(A)|^{1/p-1/q}=1$ whenever $p=q$. Since $v_{s}(Uz)\leq\sigma_{\mathrm{max}}(U)^{s}v_{s}(z)$, where $\sigma_{\mathrm{max}}(U)$ is the largest singular value of $U$, we obtain
@@
Next, we replace the window $\widehat{U}^{-1}g$ with $g$ by resorting to the standard convolution inequality in [16, Lemma 1.2.29]
@@
By Young’s inequality for weighted mixed norm spaces, see e.g. [16, Theorem 2.2.3],
(87) $\lVert|V_{g}f|\ast|V_{\widehat{U}^{-1}g}\widehat{U}^{-1}g|\rVert_{L^{p,q}_{v_{s}}}$ $\leq c^{\prime}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}}\lVert V_{\widehat{U}^{-1}g}\widehat{U}^{-1}g\rVert_{L^{1}_{v_{s}}}$
(88) $=c^{\prime}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}}\lVert V_{g}g(U\cdot)\rVert_{L^{1}_{v_{s}}}$
(89) $\leq c^{\prime}\sigma_{\mathrm{max}}(U^{-1})^{s}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}}\lVert V_{g}g\rVert_{L^{1}_{v_{s}}}$
(90) $=c^{\prime}\sigma_{\mathrm{max}}(U)^{s}\lVert V_{g}g\rVert_{L^{1}_{v_{s}}}\lVert V_{g}f\rVert_{L^{p,q}_{v_{s}}},$